A teacher surveys her class of 27 students to learn how many enjoy playing the video game Mario Kart 7 (MK7) and its newer version Mario Kart 8 (MK8). She finds that 7 students play only version 7 and that 5 students play only the newer version. If twice as many students play both versions as those that play neither version, how many students play Mario Kart 7?
Answer by macston(5194)

27 students total, 7 play only MK7 and 5 play only MK8.
27-7-5=15 There are 15 students who play neither or both.
Let n=number who play neither, 2n=number who play both.
n+2n=15
3n=15
n=5 5 students play neither
2n=2(5)=10 10 students play both.
The students who play only MK7=7 and 10 students who play both make a total of 17 students who play Mario Kart 7. ANSWER: 17